### v1

# Quantum Doubly Stochastic Transformers

Abstract
At the core of the Transformer, the softmax normalizes the attention matrix to be right stochastic. Previous research has shown that this often de-stabilizes training and that enforcing the attention matrix to be doubly stochastic (through Sinkhorn’s algorithm) consistently improves performance across different tasks, domains and Transformer flavors. However, Sinkhorn’s algorithm is iterative, approximative, non-parametric and thus inflexible w.r.t. the obtained doubly stochastic matrix (DSM). Recently, it has been proven that DSMs can be obtained with a parametric quantum circuit, yielding a novel quantum inductive bias for DSMs with no known classical analogue. Motivated by this, we demonstrate the feasibility of a hybrid classical-quantum doubly stochastic Transformer (QDSFormer) that replaces the softmax in the self-attention layer with a variational quantum circuit. We study the expressive power of the circuit and find that it yields more diverse DSMs that better preserve information than classical operators. Across multiple small-scale object recognition tasks, we find that our QDSFormer consistently surpasses both a standard Vision Transformer and other doubly stochastic Transformers. Beyond the established Sinkformer, this comparison includes a novel quantum-inspired doubly stochastic Transformer (based on QR decomposition) that can be of independent interest. The QDSFormer also shows improved training stability and lower performance variation suggesting that it may mitigate the notoriously unstable training of ViTs on small-scale data.
1 Introduction
The Transformer [52] continues to be a dominant building block powering models in natural language processing [17], computer vision [16, 32], chemistry [8] and biology [3]. At the same time, quantum computing (QC) is a novel paradigm with the potential to become practically useful in ML [43, 23, 34, 25, 1] and fuel applications across disciplines [5, 14]. Many attempts have been made to build Transformers with quantum gates, either entirely [29, 39] or only the attention blocks [28, 19, 51]. However, rather than merely migrating, recent work in quantum ML identified constraints in specific flavors of neural networks (NN) and successfully mitigated those through quantum – e.g., fourier NNs [63], graph NNs [50] or input-convex NNs [37]. Some known limitations of Transformers are due to the softmax in the attention block, e.g., entropy collapse [62], rank collapse [40], token uniformity [15], eureka moments [24] and more [59, 11, 45, 12, 61]. Applying softmax enforces the attention matrices to be right-stochastic (i.e., rows sum to $1$). Its temperature controls the entropy of this distribution, which is often adjusted to stabilize training [24, 40].
Concurrently, Sander et al. [42] discovered that Transformer attention naturally converge to doubly stochastic matrices (DSMs) over training, i.e. their rows and columns sum to 1. Motivated by this, they discovered that enforcing bistochasticity boosts Transformer performance across different modalities (text, images, point clouds). Intuitively, doubly stochastic attention has a similar effect to increasing temperature (entropy) – attention becomes more ”democratic”, less interactions are missed and all tokens are being attended more equally. Their Sinkformer [42] is a generalization of Transformers that leverages Sinkhorn’s algorithm (SA) and has been widely adopted and extended [21, 60, 44]. Among various techniques to obtain DSMs [46, 41, 13, 54], SA is the most obvious choice, however it has some disadvantages. (1) It is an iterative approximation procedure which reaches a DSM only in the limit. It is thus empirical how many iterations a Sinkformer needs. (2) It can guarantee to find a DSMs only if the input matrix is non-negative, which is generally not the case within a Transformer. (3) It is non-parametric. Thus, in contrast to e.g., a NN layer, it cannot be optimized regarding which DSM should be returned. Given the empirical superiority of the Sinkformer to vanilla Transformers, it is natural to study different techniques to make attention doubly stochastic.
Figure 1: Doubly Stochastic Transformers. Standard scaled dot-product attention applies a Softmax activation on the query-key matrix (top). We study different techniques to make attention doubly stochastic attention by replacing the softmax operation (bottom). Our proposed Quantum Doubly Stochastic Transformer (QDSFormer) leverages QontOT, a variational quantum circuit with high expressivity.
Strikingly, the recent work by Mariella et al. [37] proved (in a different context) that DSMs can be obtained naturally with a variational (i.e., parametric) quantum circuit, dubbed QontOT. They emphasize that there exists no classical learning (i.e., parametric) method that can produce a DSM, akin to QontOT. Here, we demonstrate that this opens the door for a hybrid quantum-classical doubly-stochastic Transformer (QDSFormer) which offers more flexibility than the Sinkformer. To that end, we extend QontOT to emit DSMs for an equally-sized matrix. The resulting quantum layer may replace the softmax within any standard (i.e., non-local, non-sparse) self-attention block. We focus on replacing the softmax inside the scaled dot-product self-attention of a Vision Transformer (ViT) for three reasons:
  1. 1.
    ViTs [16] suffer from unstable training [62, 24]
  2. 2.
    Unlike in NLP, the attention matrix has constant size, which eases the application of the quantum circuit
  3. 3.
    The attention matrix in a Transformer encoder is unconstrained (decoder attention matrices are masked).
We empirically analyse expressivity of the quantum circuit, finding that it yields more diverse DSMs than Sinkhorn’s algorithm both on synthetic and real data. It also preserves information better and induces higher entropy. We then train various flavors of doubly stochastic Transformers (see Figure 1) on more than ten object recognition datasets. In comparison to the ViT [16] and Sinkformer [42], the QDSFormer shows competitive performance, consistently surpassing both while stabilizing training.
2 Methods
2.1 Doubly Stochastic Matrices (DSMs)
We denote the $n$-dimensional vector of ones by $\bm{\mathrm{1}}_{n}$ and the $n\times n$ identity matrix as $\bm{\mathrm{I}}_{n}$. The Birkhoff polytope $\Omega_{n}\coloneqq\mathcal{N}\left(\bm{\mathrm{1}}_{n},\bm{\mathrm{1}}_{n}\right)$ [9] defines the convex set of $n\times n$ doubly stochastic matrices (DSMs). A DSM $\bm{\mathrm{P}}\in\Omega_{n}$ is a non-negative matrix with row/column sum of 1
$\bm{\mathrm{P}}\bm{\mathrm{1}}_{n}=\bm{\mathrm{1}}_{n},\quad$ $\bm{\mathrm{P}}^{\top}\bm{\mathrm{1}}_{n}=\bm{\mathrm{1}}_{n},\quad\bm{\mathrm{P}}_{i,j}\geq 0.$ (1)
A right stochastic matrix $\bm{\mathrm{R}}$ has row sums of 1
$\bm{\mathrm{R}}\bm{\mathrm{1}}_{n}=\bm{\mathrm{1}}_{n},\quad\bm{\mathrm{R}}_{i,j}\geq 0,$ (2)
and a left stochastic matrix $\bm{\mathrm{L}}$ has column sums of 1
$\bm{\mathrm{L}}^{\top}\bm{\mathrm{1}}_{n}=\bm{\mathrm{1}}_{n},\quad\bm{\mathrm{L}}_{i,j}\geq 0.$ (3)
Hence, a DSM is left and right stochastic. Moreover, the Birkhoff-von Neumann theorem states that the $n!$ vertices (i.e., extreme points) of the Birkhoff polytope $\Omega_{n}$ are permutation matrices, so their entries belong to $\{0,1\}$. Notably, every DSM $\bm{\mathrm{P}}\in\Omega_{n}$ can be decomposed as a convex combination of permutation matrices, that is $\bm{\mathrm{P}}=\sum_{i=1}^{N}\lambda_{i}\bm{\mathrm{\Pi}}_{i}$ for some probability vector $\bm{\lambda}\in\Delta_{N}$, $n\times n$ permutation matrices $\{\bm{\mathrm{\Pi}}_{i}\}$, and the number of extreme points $N\leq n^{2}$. While the decomposition is not unique, each DSM can be represented by at most $n^{2}$ permutation matrices [7]. Due to the linear equality constraints, the Birkhoff polytope $\Omega_{n}$ lies within a $(n-1)^{2}$-dimensional affine subspace of the $n^{2}$ dimensional space of $\mathbb{R}^{n\times n}$ matrices.
2.2 Attention
We study extensions of dot product attention [52]
$\text{Attention}(\bm{\mathrm{Q}},\bm{\mathrm{K}},\bm{\mathrm{V}})=\bm{\mathrm{A}}\bm{\mathrm{V}}=\text{Softmax}\left(\frac{\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}}{\tau}\right)\bm{\mathrm{V}}$ (4)
where $\bm{\mathrm{Q}}\coloneqq\bm{\mathrm{X}}\bm{\mathrm{W}}_{Q}$, $\bm{\mathrm{K}}\coloneqq\bm{\mathrm{X}}\bm{\mathrm{W}}_{K}$ and $\bm{\mathrm{V}}\coloneqq\bm{\mathrm{X}}\bm{\mathrm{W}}_{V}$ map the input $\bm{\mathrm{X}}$ to query $\bm{\mathrm{Q}}$, key $\bm{\mathrm{K}}$ and value $\bm{\mathrm{V}}$ through their respective weight matrix $\bm{\mathrm{W}}_{i}$ s.t. $\bm{\mathrm{Q}},\bm{\mathrm{K}},\bm{\mathrm{V}}\in\mathbb{R}^{T\times d}$. Note that $\tau^{-1}\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}\in\mathbb{R}^{T\times T}$, so the unnormalized attention matrix is quadratic. Applying the softmax operator, denoted $S(\bm{\mathrm{z}})_{i}=\frac{\exp(\bm{\mathrm{z}}_{i})}{\sum_{j=1}^{n}\exp(\bm{\mathrm{z}}_{j})}$, over the rows makes $\bm{\mathrm{A}}$ right-stochastic, i.e., each row $i$ contains a probability distribution denoting the amount of “attention” token $i$ pays to the other tokens. The temperature $\tau$, canonically set to $\sqrt{d_{k}}$ [52], controls the entropy of the output: low temperature yields a peaky distribution emphasizing differences whereas high temperature attenuates differences thus increasing entropy.
2.3 Doubly-Stochastic Operators
Below we define a non-exhaustive set of operators that can transform $\bm{\mathrm{M}}\in\mathbb{R_{+}}^{T\times T}$ to a DSM $\bm{\mathrm{P}}\in\Omega_{T}$. The operators can be integrated into a Transformer by $\bm{\mathrm{M}}\coloneqq\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}$ thus yielding a Doubly Stochastic Transformer (”DSFormer”).
2.3.1 Sinkhorn’s algorithm
The most natural approach to obtain a doubly stochastic Transformer was pursued in the Sinkformer [42] and leverages Sinkhorn’s algorithm [46]. Sinkhorn’s algorithm is based on Sinkhorn’s theorem, stating that for any square strictly positive matrix $\bm{\mathrm{M}}\in\mathbb{R_{+}}^{T\times T}$, there exist (strictly) positive diagonal matrices $\bm{\mathrm{P}}=\bm{\mathrm{D}}_{1}$, $\bm{\mathrm{D}}_{2}$ s.t., $\bm{\mathrm{D}}_{1}\bm{\mathrm{M}}\bm{\mathrm{D}}_{2}\in\Omega_{T}$. Sinkhorn’s algorithmis an approximation procedure that iteratively normalizes the mass of the rows and the columns of $\bm{\mathrm{M}}$ which has been proven to converge to a DSM by minimizing KL divergence [48]. The sole hyperparameter of this procedure is $K$, the number of iterations, which we enforce to be odd, following [42], to ensure the resulting matrix is at least numerically row-stochastic, like for the canonical Softmax operator. Moreover, we study and compare two implementations of Sinkhorn’s algorithm (SA), Naive and OT. Naive alternates between column- and row-normalization: at even iterations ($t$), each column is normalized as $\bm{\mathrm{P}}_{ij}^{(t+1)}=\bm{\mathrm{P}}_{ij}^{(t)}/{\sum_{i}\bm{\mathrm{P}}_{ij}^{(t)}}$, and at odd iterations, each row is normalized as $\bm{\mathrm{P}}_{ij}^{(t+1)}=\bm{\mathrm{P}}_{ij}^{(t)}/{\sum_{j}\bm{\mathrm{P}}_{ij}^{(t)}}$. Instead, the OT flavor is the operator used in the Sinkformer [42] which relies on the more robust and generalized version proposed by Cuturi [13] to compute optimal transport distances. Note that, both flavors may not converge with few iterations, especially if $\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}$ contains large numeric values. Therefore, the Sinkformer is only an approximately doubly stochastic Transformer.
2.3.2 Projection on the Birkhoff polytope
Alternatively, one can project $\bm{\mathrm{M}}$ directly on $\Omega_{T}$ via $\bm{\mathrm{P}}=\underset{\bm{\mathrm{X}}\in\Omega_{T}}{\operatorname*{arg\,min}}\|\bm{\mathrm{X}}-\bm{\mathrm{M}}\|_{F}^{2}$, where the set for $\bm{\mathrm{X}}$ and the objective are convex. We chose to minimize the Frobenius norm here but note that different distances could be explored. The resulting problem is a positive-definite convex quadratic program and can be rewritten as
$\min_{{\begin{subarray}{c}s.t.\\ \bm{\mathrm{x}}_{i}\geq 0\\ \bm{\mathrm{A}}\bm{\mathrm{x}}=\mathbf{1}_{2n}\end{subarray}}}\tfrac{1}{2}\bm{\mathrm{x}}^{\top}\bm{\mathrm{x}}-\bm{\mathrm{q}}^{\top}\bm{\mathrm{x}},\quad\bm{\mathrm{A}}={\text{\scriptsize$\begin{pmatrix}\bm{\mathrm{1}}_{n}^{\top}&\bm{\mathrm{0}}^{\top}&\ldots&\bm{\mathrm{0}}^{\top}\\ \bm{\mathrm{0}}_{n}^{\top}&\bm{\mathrm{1}}_{n}^{\top}&\ldots&\bm{\mathrm{0}}^{\top}\\ \vdots&\vdots&\ddots&\vdots\\ \bm{\mathrm{0}}^{\top}&\bm{\mathrm{0}}^{\top}&\ldots&\bm{\mathrm{1}}_{n}^{\top}\\ \bm{\mathrm{I}}_{n}&\bm{\mathrm{I}}_{n}&\ldots&\bm{\mathrm{I}}_{n}\\ \end{pmatrix}$}}$ (5)
[…]
2.3.3 QontOT
QontOT is a parameterized (variational) quantum circuit that was conceived very recently for conditional prediction of optimal transport plans [37]. Importantly, the circuit naturally emits DSMs which are rescaled with classical methods to transport plans. While [37] do not find signs of quantum advantage for their main task of optimal transport plan prediction, they report accuracy surpassing their classical baselines for the prediction of DSMs. This is likely a consequence of the choice of the ansatz which explores a previously unreported link between unitary operators and DSMs. Indeed, they first proved that DSMs can be obtained naturally with quantum computers thus constructing a quantum inductive bias for DSMs. Notably, as the authors state, it is currently unknown whether a similarly natural classical approach exists to produce DSMs parametrically.
For any unitary matrix $\bm{\mathrm{U}}$: $\bm{\mathrm{U}}\odot\overline{\bm{\mathrm{U}}}\in\Omega_{n}$. Given the circuit parameters $\theta$ (typically in the hundreds) and $p\in\mathbb{R}$, QontOT obtains a DSM via $\bm{\mathrm{U}}(p;\theta)\odot\overline{\bm{\mathrm{U}}}(p;\theta)$. Mariella et al. [37] block-decompose this matrix before the classical rescaling. A notable detriment of QontOT is the data injection which is limited to a scalar $p$. Therefore, we extend the multiplicative data injection $f(\theta,p)=p\cdot\theta$ from scalars to tensors, such that $f(\theta,\bm{\mathrm{M}})=\theta\odot\overrightarrow{\bm{\mathrm{M}}}$. If $\bm{\mathrm{M}}$ has less items than $\theta$ we repeat its values to obtain a vector of length identical to $\theta$. Furthermore, QontOT requires the DSM dimension $n$ to be a power of $2$. While this may be prohibitive within a Transformer (because sequence length $T$ may differ), it can be mitigated by padding. Padding to powers of two is a common technique to maximize hardware efficiency. Here, we focus our experiments on ViTs where $T$ is a direct function of patch size. In general, the circuit size scales favorably in $\mathcal{O}(\log_{2}(T))$, to be precise, it needs at least $4(\log_{2}(T)+1)$ qubits, equivalent to $2(q_{d}+q_{a})$ where $q_{d}$ is the number of data qubits ($\log_{2}(T)$) and $q_{a}$ is the number of auxilliary qubits ($\geq\log_{2}(T)+1$). In practice $T=8$.
2.3.4 QR Decomposition
As highlighted above, any unitary $\bm{\mathrm{U}}$ can provide a DSM by taking $\bm{\mathrm{U}}\odot\overline{\bm{\mathrm{U}}}$. For any input matrix $\bm{\mathrm{M}}$, we can obtain a unitary $\bm{\mathrm{U}}$ by computing an orthonormal basis for its column space. While there are many ways to obtain a basis, we choose a QR decomposition $\bm{\mathrm{M}}=\bm{\mathrm{U}}\bm{\mathrm{R}}$, in which case $\bm{\mathrm{R}}$ is upper triangular. [^1] We use $\bm{\mathrm{U}}$ rather than $\bm{\mathrm{Q}}$ for the basis of the QR to avoid confusion with the Transformer query matrix. QR has the advantage that, when implemented with Gram-Schmidt, it is differentiable if $\bm{\mathrm{M}}$ is full-rank. However note that for long-context applications the rank is rarely full because the query and key matrix have $d=\frac{d_{embed}}{n_{heads}}$ rows and typically $\bm{\mathrm{M}}\in\mathbb{R}^{T\times T}$ has rank $\min\{d,T\}$, implying that $\bm{\mathrm{M}}$ only has full rank when $d\geq T$. In practice, if the rank is defective, we inject additive Gaussian noise $\mathcal{N}(0,1\mathrm{e}{-7})$ to obtain full ranks. In the ViTs studied in our experiments, $\bm{\mathrm{M}}$ often has close to, or full-rank since the dimension $d$ is greater than the number of patches $P$, where $P\approx T$. Moreover, QR has time complexity $\mathcal{O}(n^{3})$ for dense $n\times n$ matrices, thus to scale up, approximation techniques may be needed [22].
3 Expressivity of Doubly-Stochastic Operators
Given the empirical superiority of the Sinkformer to the vanilla (i.e., right-stochastic) Transformers, a natural question is which operator to chose to obtain DSMs. Before training the DSFormers, we compare the expressivity of the operators – especially QontOT and Sinkhorn’s algorithm – in isolation on synthetic data. We focus on two aspects.
  1. 1.
    Soundness – does the operator always produce a DSM? Given that $\bm{\mathrm{U}}\odot\overline{\bm{\mathrm{U}}}\in\Omega_{n}$, QontOT always yields a DSM. Similarly for the QR decomposition. Instead, Sinkhorn’s algorithm (SA) may fail to produce a DSM if the input matrix is not positive. Within the Transformer where $\bm{\mathrm{M}}\coloneqq\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}$, the positivity requirement is generally not fulfilled which is mitigated by exponentiation of the inputs. Thus, following Sinkhorn’s theorem, SA always converges given enough iterations $k$. In practice however, the iterative procedure is limiting. When passing $8\times 8$ $\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}$ matrices from a trained Sinkformer, we observe that SA does not converge for the common choices of $k$ ($3$ and $21$). Indeed, the Frobenius distance to the Birkhoff polytope is $0.84_{\pm 0.26}$ for $k=3$ and $0.23_{\pm 0.16}$ for $k=21$. This is in contrast to the QR, QontOT and the Birkhoff projection which all yield distances $<2\mathrm{e}{-4}$, QontOT even $<5\mathrm{e}{-6}$. Instead, the vanilla Softmax operator yields a right-stochastic matrix with distance $1.12_{\pm 0.33}$. Hence Sinkformer attention is only approximately doubly stochastic.
  2. 2.
    Completeness – can the operator produce all possible DSMs? Sinkhorn’s algorithm reaches all DSMs of the form $\bm{\mathrm{P}}=\bm{\mathrm{D}}_{1}\bm{\mathrm{M}}\bm{\mathrm{D}}_{2}$. However, due to the entry-wise exponentiation of $\bm{\mathrm{M}}$ in the Sinkformer, the input matrix never contains any zero, thus the boundaries of the polytope cannot be reached. Regarding QontOT, the resulting DSM is a convex combination of unistochastic matrices [37][Eq. 11b]. Unistochastic matrices are a non-convex proper subset, covering a large amount of the Birkhoff polytope (albeit the exact amount is unknown [18]). In theory, if all unistochastic matrices could be reached, then by their convex combinations QontOT could cover the entire Birkhoff polyope. In practice, reaching all unistochastic matrices (especially all permutation matrices) with the same circuit parametrization is unfeasible as it requires fault-tolerant quantum hardware and high circuit depth (entanglement). But over the parameter space of QontOT, the Birkhoff polytope can be approximated more closely.
3.1 Empirical analysis
Figure 2: Number of unique DSMs obtained with each operator after exhaustively iterating over a discretized unit hypercube. With only $8$ layers, QontOT produces a unique DSM for every possible input, unlike all other methods.
To empirically assess the completeness of the operators w.r.t. the Birkhoff polytope, we performed a brute-force analysis over a discretized grid of the unit hypercube. For a $n\times n$ matrix and a discretiztion step $d\in\mathbb{N}_{+}$, we sample each column from a discretized $n$-dimensional hybercube with $d^{n}$ points, yielding $d^{n^{2}}$ unique matrices. For $n=4$ and $d=3$ we obtain $3^{16}\approx 43$M matrices and computed the DSM for each input, before rounding to third decimal place. Across all operators, QontOT yielded by far the most unique DSMs (see Figure 2), behaving nearly injective when $8$ or more circuit layers are used. This is important, because, none of the other operators is injective thus some information is lost when using it inside a neural network. A closer inspection of the empirical cumulative distribution function of all DSMs reveals that QR often emits the same DSMs and that with only $2$ layers (i.e., $98$ parameters), QontOT surpasses all other methods (see Appendix Figure A2). Furthermore, whereas all the classical techniques are non-parametric, QontOT is a learning method, thus each parameter configuration yields a different set of DSMs. In general, Sinkhorn’s algorithm and the direct Birkhoff projection are both permutation and rotation equivariant. Instead, QR and QontOT do not possesses this characteristic. Moreover, the Naive flavor of SA and the QR are scale-invariant whereas OT and QontOT are not. Note that such equivariances within a Transformer are not generally beneficial or detrimental. We also examined the operators on $8\times 8$ $\bm{\mathrm{Q}}\bm{\mathrm{K}}$ matrices from a single batch ($N=3584$, $7$ heads, batch size of $512$) of a trained Sinkformer. In general, the operators were injective on this set, yielding unique DSMs, however the OT flavor of Sinkhorn’s algorithm only produced $2175$ ($2021$) unique DSMs for $k=3$ ($21$) iterations. This deficiency may be due to the inputs having large numerical values and not being generally non-negative.
Beyond approximate injectivity, we argue that a powerful operator needs to possess two further characteristics: First, information has to be preserved. Obtaining unique DSMs is useless if they destroy information from the input matrix. To assess this, we measured the Frobenius norm of the residuals between input and output matrix [^2] Other metrics like measuring preservation of ranks or pairwise ratios are possible but yielded similar results.. Secondly, low entropy has to be avoided because it causes vanishing gradients and destabilizes Transformer training [62, 45] for which various mitigation techniques have been suggested [55, 24]. This so-called “entropy collapse” arises if attention is too spiky and is induced by low temperature in the softmax. Our analysis in Figure 3 reveals that QontOT possess better information preservation and larger entropy than Sinkhorn’s algorithm on realistic unnormalized attention matrices.
Figure 3: Shannon entropy of different flavors of doubly stochastic attention against a measure of distance preservation – the Frobenius norm of the difference between unnormalized attention $\bm{\mathrm{Q}}\bm{\mathrm{K}}$ and their obtained DSM $\bm{\mathrm{P}}$. QontOT uses $2$ circuit layers.
Among all operators, QontOT possess the best information preservation properties while still maintaining Shannon entropy that is only surpassed by the QR decomposition.
Next, we assessed which combination of circuit layers and auxiliary qubits yields the best speed-expressivity compromise. In general, a single circuit execution is in the three-digit millisecond range but can be efficiently parallelized. Increasing the number of layers causes a sub-linear runtime increase, whereas adding more qubits causes a exponential increase (Appendix Figure A1). Regarding expressivity, adding more layers has a higher impact than adding more qubits (Appendix Figure A3). In detail, we passed the same matrix $10,000$ times, sampled the circuit parameters $\theta\sim\mathcal{U}(-1,1)$ and then measured the average range of values covered within each cell of the DSM. This shows that adding more auxiliary qubits is only useful if even more layers are added simultaneously.
Altogether, these results are promising and motivate our below empirical experiments with a Quantum Doubly Stochastic Transformer (QDSFormer).
4 Theoretical result on number of DSMs
The optimal way of studying the expressivity of a doubly-stochastic operator empirically would enumerate all DSMs in a given Birkhoff polytope $\Omega_{n}$ and assess for each DSM whether it can be reached (or how closely). The exact volume of the Birkhoff polytope is an open problem in mathematics [10] which limits our ability to study expressivity theoretically. In practice, one can assume a certain discretization $p\in\mathbb{N}_{+}\text{ s.t., }\bm{\mathrm{P}}_{ij}\in\{0,\frac{1}{p-1},...,1\}$, e.g., if $p=2$ then $\bm{\mathrm{P}}_{ij}\in\{0,1\}$. In that case there are $n!$ DSMs. In Appendix D we provide a partial derivation for the combinatorial problem of identifying the function $f(n,p)\rightarrow\mathbb{N}$ returning the number of DSMs. The basic idea is as follows. Given that a $n\times n$ DSM has $(n-1)^{2}$ degrees of freedom, there are $p^{(n-1)^{2}}$ candidate matrices. Not all of these can be turned into DSMs because of two constaints, (1) the sum of any row or column must not exceed 1 and (2) the sum of the $n-1\times n-1$ submatrix must not be below $n-2$ [10]. This allows to decompose $f$ into $f(n,p)=p^{(n-1)^{2}}-c_{1}-c_{2}+c_{12}$ where $c_{1}$ and $c_{2}$ measure how often the constraints are violated and $c_{12}$ corrects for the cases where both constraints are violated. For details see Appendix D.
5 Quantum Doubly Stochastic Transformer
5.1 Experimental Setup
We evaluate different flavors of DSFormers obtained through replacing the Softmax function with any of the DSM operators described above. When integrating our QontOT-derived operator into a ViT we obtain our hybrid quantum-classical doubly stochastic Transformer (QDSFormer). In the following, we refer to QontOT as the attention flavor which contains the quantum circuit whereas QDSFormer denotes, more broadly, any Transformer with quantum doubly-stochastic attention. To date, the only realization of a QDSFormer is through QontOT. Among all operators, the classical ones are non-parametric whereas this quantum operator can be optimized during training. Therefore, one could theoretically optimize circuit parameters concurrently with Transformer training. However, the ViTs we study contains up to $4$ attention layers, with a batch size of $512$, yielding $2048$ samples to optimize in a single forward pass. We predict, unless mentioned otherwise, $8\times 8$ DSMs, use $16$ circuit layers and $4$ auxiliary qubits ($16$ qubits in total). Running the circuit on quantum hardware requires $\Omega(n^{2}/\varepsilon^{2})$ shots to obtain satisfactory sampling error [37]. Assuming a precision of $\varepsilon=0.01$, this is in the order of $640$k shots per sample. Since quantum hardware operates on kHz frequency, execution and online optimization on hardware is unfortunately not (yet) feasible. Therefore, we perform exact statevector simulation with Qiskit [26] and implement three circuit training strategies:
  1. 1.
    Differentiable: Joint optimization of circuit and Transformer parameters through backpropagation, akin to integrating the circuit as a neural network layer. This is by far the slowest given the difficulty of gradient propagation through quantum circuits [2].
  2. 2.
    Mixed: A mixed strategy where Transformer training is interleaved with $200$ steps of gradient-free circuit optimization with Nevergrad [6] on a per-epoch basis.
  3. 3.
    Static: The circuit is used in pure inference mode with parameters obtained from a 24-qubit DSM prediction experiment on quantum hardware [37].
From the operators studied in Section 3.1, we discard the Birkhoff projection due to its non-differentiability and low DSM-diversity (Figure 2). For comparison, we further include the NormSoftmax [27], here denoted as Softmaxσ, that attenuates attention by taking the minimum of the empirical and the expected standard deviation $\tau:=\sqrt{d_{k}}$:
$\bm{\mathrm{A}}=\text{Softmax}\left(\frac{\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}}{\min{\left(\sigma(\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}),\tau\right)}}\right).$ (6)
[…]
5.2 Data sets
We evaluate the different ViTs on MNIST [33], Fashion MNIST [57], seven datasts from the MedMNIST benchmark [58] and a compositional task requiring multistep reasoning as proposed by Hoffmann et al. [24]. In that task, a $2\times 2$ grid contains two MNIST digits (upper left and lower right) and two FashionMNIST items (upper right and lower left). If the digits have equal value, the label is the upper right fashion item, otherwise it is the bottom left fashion item. Performance typically ramps up quickly to $\sim 50\%$ because the model learns to attend one (and only one) of the FashionMNIST images. Upon continued training with a long saturation phase, a ViT suddenly grasps the relationship of the MNIST digits to the classification task and then climbs rapidly to a $90-95\%$ accuracy. The moment of abrupt improvement is called ”Eureka moment” [24]. The dataset is split into $60$K ($10$K) training (validation) examples. To accommodate the $8\times 8$ attention matrix, each image from MNIST, FashionMNIST and MedMNIST is split into $7$ horizontal stripes and a CLS token is pre-pended to the patch sequence. For the Eureka dataset we use a patch size of $14\times 28$ pixels and mean-pooling.
5.3 Empirical results
First, we compare the QDSFormer directly with a standard ViT. The ViT uses softmaxed attention whereas the QDSFormer employs a ViT with a static (i.e., non-trainable) instantiation of QontOT to make attention doubly stochastic. Figure 4 clearly indicates that on both datasets, the QDSFormer exceeded the ViT by a significant margin.
(a) 1-ViT-Layer on MNIST.
(b) 2-ViT-Layer on FashionMNIST.
Figure 4: Comparison of ViT and QDSFormer while varying the circuit depth. Mean/std from 5 trainings are shown. Within (a) and (b) all models use the same number of trainable parameters.
This confirms the finding that doubly stochastic attention can improve ViTs [42]. Moreover, in both cases, adding more circuit layers increases performance logarithmically and with $4$ or $8$ circuit layers the ViT performance is surpassed. Exact numerical results are provided in Appendix Table A3.
[…]
Table 1: $L$-layered ViT validation accuracy on FashionMNIST for different attention methods. QontOT uses $16$ circuit layers. Mean/std computed from $5$ trainings.
[…]
1 $\underline{86.5}_{\pm 0.19}$ $75.3_{\pm 4.61}$ $\textbf{87.1}_{\pm 0.26}$ $85.6_{\pm 0.10}$ $84.2_{\pm 3.64}$
2 $88.9_{\pm 0.12}$ $84.6_{\pm 2.11}$ $\underline{89.3}_{\pm 0.07}$ $\textbf{90.0}_{\pm 0.15}$ $89.1_{\pm 0.73}$
3 $\underline{89.4}_{\pm 0.28}$ $86.3_{\pm 2.69}$ $\underline{89.4}_{\pm 0.11}$ $\textbf{90.3}_{\pm 0.13}$ $\underline{89.4}_{\pm 0.77}$
4 $\underline{89.7}_{\pm 0.29}$ $87.1_{\pm 1.15}$ $89.5_{\pm 0.07}$ $\textbf{90.3}_{\pm 0.14}$ $89.1_{\pm 1.08}$
The same result was obtained on MNIST and, this time, QontOT outperformed softmaxed attention also for one ViT layer (see Table 2).
Table 2: Identical to Table 1 but for MNIST.
[…]
1 $89.1_{\pm 12.5}$ $66.7_{\pm 22.5}$ $\textbf{96.6}_{\pm 0.10}$ $93.9_{\pm 0.11}$ $\underline{94.3}_{\pm 1.97}$
2 $98.1_{\pm 0.33}$ $93.0_{\pm 4.57}$ $\underline{98.3}_{\pm 0.13}$ $\textbf{98.4}_{\pm 0.05}$ $98.2_{\pm 0.27}$
3 $\underline{98.6}_{\pm 0.11}$ $97.7_{\pm 0.65}$ $\underline{98.6}_{\pm 0.13}$ $\textbf{98.7}_{\pm 0.06}$ $\underline{98.6}_{\pm 0.12}$
4 $\textbf{98.8}_{\pm 0.10}$ $97.9_{\pm 0.71}$ $\underline{98.7}_{\pm 0.11}$ $\textbf{98.8}_{\pm 0.07}$ $97.9_{\pm 1.57}$
In further experiments with more ViT layers performance assimilated and plateaued due to the simplicity of the datasets. But we saw scant further improvement for more than $16$ circuit layers. For a barplot visualization of Table 1 and 2 see Appendix Figure A4/A5. Notably, QontOT offers great flexibility in the type of ansatz for the quantum circuit [37]. We observed only minor differences between four different ansatz types, with three of them outperforming the ViT, underlining the generality of the finding (Appendix Table A1).
A compelling finding is that the static version of the QontOT-attention did perform as good or even better than the optimized one (see Appendix Figure A7). We tested an end-to-end optimizable QDSFormer where circuit and ViT parameters are jointly optimized. Such end-to-end training is not only slower, but also had lower accuracy than the static configuration, for both MNIST and FashionMNIST and $1$ and $8$ circuit layers (Figure A7). This may be caused by Barren plateaus [38] (i.e., gradients are largely constant along most directions), a widespread phenomenon in variational quantum circuits that slows down learning. We further experimented with a ”mixed” training strategy where the circuit is trained every $n$-th epoch. This did not reveal a clear benefit for more frequent circuit optimization (see Appendix Figure A6), potentially due to higher volatility of the circuit. We therefore use the static, faster configuration in all remaining experiments.
Table 3: Test accuracy for different MedMNIST datasets across five attention types in a 2-layer ViT. QontOT uses 16 circuit layers.
[…]
dataset
OCT $\textbf{64.4}_{\pm 1.6}$ $43.6_{\pm 3.0}$ $62.5_{\pm 0.9}$ $61.6_{\pm 0.6}$ $55.1_{\pm 5.2}$
Pneumonia $84.2_{\pm 0.8}$ $84.7_{\pm 2.0}$ $84.3_{\pm 0.7}$ $\textbf{86.1}_{\pm 1.0}$ $83.0_{\pm 1.5}$
Tissue $60.0_{\pm 0.2}$ $49.4_{\pm 1.2}$ $59.0_{\pm 0.1}$ $\textbf{60.6}_{\pm 0.1}$ $56.9_{\pm 2.0}$
OrganA $78.8_{\pm 0.5}$ $73.6_{\pm 1.7}$ $78.4_{\pm 0.6}$ $\textbf{81.2}_{\pm 0.3}$ $77.0_{\pm 2.5}$
OrganC $79.8_{\pm 0.5}$ $71.7_{\pm 7.3}$ $79.6_{\pm 0.3}$ $\textbf{82.7}_{\pm 0.5}$ $79.7_{\pm 1.0}$
OrganS $64.4_{\pm 0.6}$ $59.3_{\pm 0.9}$ $62.6_{\pm 0.8}$ $\textbf{68.1}_{\pm 0.6}$ $63.5_{\pm 0.9}$
Breast $79.6_{\pm 2.0}$ $78.2_{\pm 2.2}$ $\textbf{81.3}_{\pm 2.9}$ $80.0_{\pm 1.1}$ $80.1_{\pm 0.8}$
Mean $73.0$ $65.8$ $72.5$ 74.3 $70.8$
Next, we repeated the comparison to the four classical attention types on larger datasets (up to $240$k images) from the MedMNIST benchmark [58]. In $5$ out of $7$ datasets, the QDSFormer obtained significantly better results than all other methods (Table 3), with a mean accuracy increase of $1.3\%$ compared to a standard ViT. Notably, none of the other attention types can improve upon the standard ViT.
Another important advantage of the QDSFormer is its stabilizing effect. Among repeated training runs, the performance variation (i.e., test accuracy variance) is consistently lower than for all classical methods (e.g., Table 2 or Table 3). Notably, with 1 ViT-layer and softmax-based attention some trainings on MNIST failed to converge.
[…]
Figure 5: Validation accuracy for each training epoch, highlighting the abrupt learning referred to as Eureka Moment (EM) on the compositional dataset. Confidence bounds from 5 runs.
This major improvement, achieved with an extremely lightweight quantum circuit (1 layer) also consolidated across different learning rates (Table 4).
Table 4: Results on the Eureka dataset across different models and learning rates. EM@Ep denotes the average epoch of the EM, for runs without EM, the epoch is considered to be $100$.
[…]
$1$e-$3$ Acc. $53.4_{\pm 0.09}$ $48.6_{\pm 1.12}$ $49.0_{\pm 0.85}$ $\underline{61.2}_{\pm 11.1}$ $\textbf{70.0}_{\pm 13.9}$
EM@Ep $\textbf{72.2}_{\pm 14.2}$ $\underline{74.8}_{\pm 8.43}$
# EM $0/5$ $0/5$ $0/5$ 2/5 2/5
$7$e-$4$ Acc. $53.9_{\pm 0.16}$ $50.9_{\pm 0.83}$ $50.9_{\pm 0.65}$ $\underline{72.9}_{\pm 10.6}$ $\textbf{82.3}_{\pm 14.4}$
EM@Ep $\underline{57.7}_{\pm 10.6}$ $\textbf{43.3}_{\pm 7.01}$
# EM $0/5$ $0/5$ $0/5$ 4/5 4/5
$5$e-$4$ Acc. $61.1_{\pm 14.6}$ $51.0_{\pm 0.52}$ $51.6_{\pm 0.24}$ $\underline{66.4}_{\pm 15.6}$ $\textbf{89.4}_{\pm 0.09}$
EM@Ep $61.0$ $\underline{72.6}_{\pm 21.6}$ $\textbf{43.6}_{\pm 8.25}$
# EM $1/5$ $0/5$ $0/5$ $\underline{\vphantom{y}\smash{2/5}}$ 5/5
Next, to assess the potential of the QDSFormer beyond image data and simultaneously study the scalability of the QDSFormer, we applied it on a novel dataset for time-series classification of InfraRed (IR) spectra of molecules into $37$ functional groups [4]. This dataset contains almost $1$M samples and we scale up the circuit to produce $4$x larger attention matrices ($16\times 16$). On this dataset, the performance differences are marginal and the QDSFormer performs on par with a standard ViT (for details se Appendix Table A2). This shows that the QDSFormer can meaningfully generalize to domains like scientific data and tasks like multi-label classification.
Since all above results were obtained through statevector simulation, we conducted a final experiment to understand the detrimental effect of quantum noise induced by real quantum hardware (via the publicly available IBM Quantum Platform, for details see Appendix E). This 14-qubit experiment tests the potential for a hybrid hardware training. Despite various light error mitigation techniques, the obtained doubly stochastic attention matrices consistently show high entropy (i.e., a tendency toward more uniform distributions), even for larger shot counts.
Figure 6: Hardware experiment on different quantum computers available via IBM Quantum Platform. A The Frobenius distance between the hardware-obtained DSM to its noise-free equivalent. B: The spearman rank correlation between the $64$ values in the noise-free and hardware-obtained DSMs show that the ordering of values is largely preserved. D: Execution time in seconds on different quantum computers.
However, as Figure 6B shows, when comparing to the noise-free ground truth attention matrix, the ordering of the values in the attention matrix was preserved with high precision (spearman $\rho>0.9$ even for moderate shot count). Since the circuit runs within a ViT, successfully preserving the ordering will be key (to not destroy signal). Instead, numerical exactness (cf. Figure 6A) may be compromised and we suspect that embedding such a noisy quantum attention block (which preserves peak attention scores but also increases entropy) into a Transformer could even be advantageous. The additional entropy may avoid vanishing gradients and act as a form of regularization. This effect is particularly notable compared to its noise-free analog, which remains classically intractable if sufficient qubits are used.
6 Conclusion
Here, we proposed the quantum doubly stochastic Transformer. We conceived this method by connecting the centerpiece of a novel variational quantum circuit (devised for optimal transport [37]) with the Transformer, facilitated through the empirical observation by Sander et al. [42] that doubly-stochastic attention improves performance in Transformers. By extending the QontOT circuit from scalars to matrices we enabled its integration into a ViT, thus providing the first parametric, doubly-stochastic Transformer. Notably, the QDSFormer presents a meta-class of Sinkformers because it estimates DSMs parametrically, i.e., it can be optimized to learn arbitrary transformations onto the Birkhoff polytope. Moreover, to our knowledge, there is no classical, parametric approach to estimate DSMs, thus the QDSFormer is a promising candidate for hybrid quantum-classical neural networks trained on quantum hardware.
Our empirical expressivity analysis revealed that the quantum circuit produces DSMs that are more diverse, preserve information better and have higher entropy than DSMs from Sinkhorn’s algorithm. On multiple simple object recognition tasks, the QDSFormer exhibited significantly higher accuracy, outperforming a ViT and a Sinkformer in most cases. Our usage of quantum attention substantially stabilizes the notoriously unstable ViT training on small-scale data, as evidenced by the performance on a compositional object recognition task (Figure 5), previously used to study ViT training dynamics [24]. Albeit these results are promising, all experiments were performed on comparably small-scale, due to the currently poor scaling of quantum computers in general (which is expected to improve). Notably, by leveraging QR decomposition, we also proposed a novel, quantum-inspired attention flavor. Broadly speaking, outsourcing the activation function to a parametric quantum circuit might be seen a computational overhead, however, we envision that this may reveal potential benefits (typically in small-data, small-model and short-training settings [1]) that are out of reach for classical hardware. To that end, future work could explore concurrent optimization of ViT and circuit parameters via the parameter-shift rule on real quantum hardware.
Appendix A Circuit execution times
In Figure A1 we report detailed runtimes for the QontOT algorithm for different combinations of circuit layers and auxiliary qubits.
Figure A1: QontOT circuit execution times for DSM of size $4$ (A) and $8$ (B) for different combinations of qubits and circuit layers. C and D show the relative increase in execution time as a function of increasing the number of layers (C), and qubits (D). Adding more layers has a sublinear effect on runtime, adding qubits requires exponential more runtime. The minimal number of auxiliary qubits is $\log_{2}(n)+1$ and the total number of qubits is $2(q_{d}+q_{a})$ where $q_{d}$ and $q_{a}$ are data and auxilliary qubits respectively.
Appendix B Empirical circuit expressivity
Figure A2: DSM counts were ranked descendingly and plotted against their cumulative count. QontOT generally produces more diverse DSMs than Sinkhorn’s algorithm.
Figure A3: Mean range of observed values in the DSM obtained from a single, random input matrix, when randomly sampling $1000$ circuit parametrizations.
Appendix C QDSFormer results
C.1 Hyperparameters
For experiments on MNIST, FashionMNIST and the seven MedMNIST datasets, the ViT was configured with a hidden dimension of 128 and an MLP dimension expansion factor of 1. The model was tested with 1 to 4 Transformer layers, each containing a single attention head. No dropout was applied, and the batch size was set to 100. For the optimizer, Adam [31] was used, and the learning rate schedule followed the setup in Sander et al. [42], with an initial learning rate of 5e-4, decreasing by a factor of 10 at epochs 31 and 45. For the more complex Eureka dataset, comprised of 56x56 RGB images, the hidden dimension was increased to 256, and a larger batch size of 512 was used. The MLP expansion factor was also doubled to 2. A cosine learning rate schedule was used with the optimizer AdamW [35]. The scheduler uses 5 warmup epochs with a warm-up learning rate of 1e-6, the decay rate is set to 0.1 and the minimum learning rate is 1e-5, the other parameters follows the default TIMM settings [56]. For the optimizer the weight decay is 0.05 and betas $(0.9,0.999)$. For the Eureka dataset, no Exponential Moving Average (EMA) is used.
C.2 QontOT ansatz types
Table A1: Ablation study for a $2$-layer QDSFormer with different circuit ansatz types and varying number of layers on FashionMNIST. Mean/std of $5$ runs.
Circuit L. Simple Parted Centrosymmetric Trotter
1 $\mathbf{88.0}_{\pm 0.10}$ $87.7_{\pm 0.22}$ $86.4_{\pm 0.23}$ $87.7_{\pm 0.08}$
8 $\mathbf{89.9}_{\pm 0.15}$ $89.8_{\pm 0.15}$ $89.4_{\pm 0.17}$ $88.4_{\pm 0.20}$
Simple: This ansatz is the most generic and resembles a checkerboard structure formed by 4-parameter unit-blocks acting on two qubits each [30, 36]. If all parameters are zero, it falls back to the identity. This ansatz is convenient because it is shallow in simulation but whose depth may vary depending on qubit layout of the quantum hardware.
Parted: This ansatz partitions the Simple ansatz into two parts: $U=U_{1}\otimes U_{2}$, where $U_{2}$ operates normally, and $U_{1}$ is transposed and placed around the initial Bell state. This design reduces the original Simple ansatz circuit depth nearly by half, which may be more efficient on certain quantum hardware. However depending on the qubit layout of the quantum hardware, it carries the potential of increased transpiled circuit depth, as the 2-qubit gates may act on distant qubits necessitating additional swap gates upon transpilation, which we observed on IBM Eagle and Heron quantum processing units. Unless mentioned otherwise, we used this ansatz in all our experiments as it yields shallower circuits in simulation and the increased depth compared to the Simple ansatz was negligible at tested system sizes.
Centrosymmetric: This was the predominantly used ansatz by Mariella et al. [37]. It is less generic, biasing toward properties of centrosymmetric matrices.
Trotter: This ansatz implements a second-order Trotter decomposition [47]. Each circuit layer corresponds to a Trotter step.
C.3 Time series classification
Table A2: Micro-F1 on IR spectra dataset across $5$-fold cross-validation with a 1-layer ViT. QDSFormer uses 16 circuit layers for both DSM sizes.
[…]
$8\times 8$ $81.60_{\pm 0.34}$ $81.41_{\pm 0.23}$ $\underline{81.68}_{\pm 0.18}$ $\textbf{81.70}_{\pm 0.05}$ $81.38_{\pm 0.22}$
$16\times 16$ $\textbf{81.55}_{\pm 0.07}$ $80.94_{\pm 0.13}$ $\underline{81.48}_{\pm 0.10}$ $81.06_{\pm 0.27}$ $80.98_{\pm 0.30}$
C.4 Ablation studies
Table A3: QDSFormer ablation varying the circuit layers. Exact numbers corresponding to Figure 4.
Configuration Validation Accuracy (%)
MNIST FashionMNIST
QDSFormer-1L $83.4_{\pm 0.73}$ $87.7_{\pm 0.22}$
QDSFormer-2L $85.7_{\pm 0.60}$ $88.5_{\pm 0.27}$
QDSFormer-4L $87.7_{\pm 0.73}$ $89.3_{\pm 0.18}$
QDSFormer-8L $91.8_{\pm 0.57}$ $89.8_{\pm 0.15}$
QDSFormer-16L $93.8_{\pm 0.10}$ $\textbf{90.0}_{\pm 0.15}$
QDSFormer-32L $\textbf{94.2}_{\pm 0.30}$ $\textbf{90.0}_{\pm 0.13}$
Baseline
ViT $92.9_{\pm 3.76}$ $88.9_{\pm 0.12}$
Figure A4: FashionMNIST results of different ViT layers for different attention types.
Figure A5: MNIST results of different ViT layers for different attention types.
Figure A6: Ablation on different number of QontOT trainings on MNIST in the mixed circuit optimization strategy.
C.5 Differentiable circuit
(a) QDSFormer 8 circuit layers.
(b) QDSFormer 1 circuit layers.
Figure A7: Differentiable QontOT training versus the static circuit.
Appendix D Counting DSMs
In Section D.1 we provide a full analytical solution to count the number of $3\times 3$ DSMs for a given discretization $p\in\mathbb{N}$. In Section D.2 we strive to extend the analytic solution to an arbitrary $n\in\mathbb{N}$ but only provide a partial solution. Finally, in Section D.3 we provide numerical results from brute-force calculation of the number of DSMs that verified the explicit analytical solution in Section D.1.
D.1 Analytical solution for $n=3$
Intuition. By systematically testing all combinations from the discretized range of values, starting with an initial $2\times 2$ zero matrix, each element is incrementally increased in the order $x_{0,0},x_{1,0},x_{0,1},x_{1,1}$ with the next highest value in the discretized range. Once an element reaches its maximum value, the next element is increased, and the preceding elements are reset to 0. This cycle repeats, starting again with the first element.
Explanation. Assume that $n=3$ and a specific discretization $p\in\mathbb{N}_{+}$ are given. In this scenario, the corresponding $3\times 3$ matrix possesses $4$ degrees of freedom, implying that the associated submatrix has dimensions $2\times 2$. The first constraint requires that the sum of the elements in each row and each column of the matrix must not exceed 1.
If a specific element $e_{ij}$ with $i,j\in\{0,1\}$ is chosen and assigned a value $x_{i}$, the possible values for the remaining elements in the same row and column can be determined.
Given that each element can assume exactly $p$ distinct values, the total number of combinations is computed as a sum over all $p$ values:
$f(3,p)=\sum_{i=1}^{p}c_{i}$ (7)
The possible values for the elements in the same row and column are restricted to the subset $\{x_{1},\dots,x_{p-i+1}\}$. As a result, the amount of submatrices that satisfy the first constraint can be expressed as:
$f(3,p)=\sum_{i=1}^{p}\left[\sum_{j=1}^{p-i+1}\left[\sum_{k=1}^{p-i+1}c_{ijk}\right]\right]$ (8)
To determine the possible values for the last element, it is necessary to consider the elements $e_{ij}^{\prime}$. The minimum number of possible values derived from these elements defines the number of candidates for the last element:
$f(3,p)=\sum_{i=1}^{p}\sum_{j=1}^{p-i+1}\sum_{k=1}^{p-i+1}\sum_{l=1}^{\min(p-j+1,p-k+1)}\mathbbm{1}(i,j,k,l,p)$ (9)
Up to this point, only the first constraint has been considered. To fully satisfy the problem requirements, matrices that violate the second constraint must be excluded. The second constraint is satisfied when the sum of the indices of all elements does not exceed $p$. Instead of subtracting $1-\mathbb{I}(i,j,k,l,p)$, the condition is captured using an indicator function $\mathbb{I}(i,j,k,l,p)$, defined as:
$\mathbbm{1}(i,j,k,l,p)=\begin{cases}1&\text{if }i+j+k+l-3\geq p,\\ 0&\text{otherwise}.\end{cases}$ (10)
By incorporating $\mathbbm{1}(i,j,k,l,p)$, the expression extends to:
$f(3,p)=\sum_{i=1}^{p}\sum_{j=1}^{p-i+1}\sum_{k=1}^{p-i+1}\sum_{l=1}^{\min(p-j+1,p-k+1)}\mathbbm{1}(i,j,k,l,p)$ (11)
This can be summarized as follows:
$f(3,p)=\sum_{(i,j,k,l)\in D(p)}\mathbbm{1}(i,j,k,l,p)$ (12)
where $D(p)=\{(i,j,k,l)\mid 1\leq i\leq p,1\leq j\leq p-i+1,1\leq k\leq p-i+1,1\leq l\leq\min(p-j+1,p-k+1)\}$.
This equation has been validated computationally for values up to $p=43$, demonstrating alignment with our empirical results.
D.2 General approximation
To determine the number of unique DSMs for a given $n,p\in\mathbb{N}$ we try to solve
$f(n,p)=p^{(n-1)^{2}}-c_{1}-c_{2}+c_{12}$ (13)
where the first term calculates the number of DSM-candidate matrices, $c_{1}$ and $c_{2}$ measure how often the constraints are violated and $c_{12}$ is a small correction term counting cases where both constraints are violated.
Generally, $c_{12}$ is very small, yet difficult to compute, thus a tight lower bound can be given with the remaining three terms. Below, we provide a derivation for $c_{2}$. We leave the derivation of $c_{1}$ to future work.
D.2.1 Constraint 2
Constraint 2.
The sum of the $n-1\times n-1$ inner matrix must not be below $n-2$ [10].
We aim to find a function $c_{2}(n,p)$ that computes the number of violations to Constraint 2 for a given $n,p\in\mathbb{N}$ when exhaustively looping over all $p^{(n-1)^{2}}$ candidate matrices that uniquely determine a $n\times n$ DSM.
An $n-1\times n-1$ matrix where each cell $x_{ij}$ can take $p$ values has
$|u|=(n-1)^{2}(p-1)+1$ (14)
unique possible sums. These sums are regularly spaced from 00 to $(n-1)^{2}$ with a step size of $p-1$, i.e., $u_{i}=\left\{\frac{i}{p-1}\mid i\in\{0,1,...,|u|\}\right\}$. This allows conversion to an integer problem (by multiplication of $p-1$) and apply Stars & Bars Theorem 2.
Theorem D.1.
For any $s,k\in\mathbb{N}$, the number of $k$-tuples $(x_{0},...,x_{k})$ where $x_{k}\in\mathbb{N}_{0}$ with sum $s$ is equal to the number of multisets of cardinality $s$ taken from a set of size $k$:
$\binom{s+k-1}{k-1}$ (15)
Specifically, we set $k:=(n-1)^{2}$ and then define the set of sums that violate the constraint as $S:=\{s\in\mathbb{N}_{0}\mid 0\leq s\leq(n-2)(p-1)\}$. Thus $|S|=(n-2)(p-1)$ We then compute the violations via:
$\hat{c}_{2}(n,p)=\sum_{s=0}^{(n-2)(p-1)}\binom{s+(n-1)^{2}-1}{(n-1)^{2}-1}$ (16)
Unfortunately, this is only approximately correct because Theorem D.1 assumes $x_{k}\in\mathbb{N}_{0}$, instead we require $x_{i}\in\{0,1,\dots,p-1\}$. Therefore, we exclude solutions where any $x_{i}>p-1$ through the inclusion exclusion principle as follows.
Assume that some $x_{i}>p-1$. We set $\hat{x}_{i}=x_{i}-(p-1+1)$. Since $\hat{x}_{i}\geq 0$, we can rewrite the original sum
$s=\sum x_{i}=\sum_{i\in I}\left(\hat{x_{i}}+p\right)+\sum_{i\notin I}x_{i}$ (17)
where $I$ is the set of indices where $x_{i}>p-1$.
Now let $m:=|I|$ be the number of violating variables, then
$s-mp=\sum_{i\in I}\hat{x_{i}}+\sum_{i\notin I}x_{i}$ (18)
To find the number of non-negative integer solutions to Equation 18, we can again leverage Theorem D.1, but now in a corrected form:
$d(n,p,m,s)=\begin{cases}\binom{s-mp+(n-1)^{2}-1}{(n-1)^{2}-1}&\text{if }s-mp>0,\\ 0&\text{otherwise}.\end{cases}$ (19)
This gives a solution for a specific sum $s\in S$ and number of violations $m$. However, since $0\leq m\leq(n-1)^{2}$, we have to sum over all options of $m$ and apply the inclusion-exclusion principle to avoid over-/undercounting.
Figure A8: Number of DSMs of fixed size $n$ with a given number of discretization steps $p$; up to values of $n=4$ and $p=30$.
$c_{2}(n,p,s)=\sum_{m=0}^{(n-1)^{2}}\left[(-1)^{m}\binom{(n-1)^{2}}{m}d(n,p,m,s)\right]$ (20)
where $\binom{(n-1)^{2}}{m}$ accounts for the number of ways to choose $m$ out of $(n-1)^{2}$ variables that exceed $p-1$.
Plugging Equation 20 back into the initial summation over all values $s\in S$ violating the constraint (see Equation 16), we obtain the final formula:
$c2(n,p)=\sum_{s=0}^{(n-2)(p-1)}c_{2}(n,p,s)\\ =\sum_{s=0}^{|S|}\sum_{m=0}^{(n-1)^{2}}(-1)^{m}\binom{(n-1)^{2}}{m}\begin{cases}0&\text{if }s-mp\leq 0,\\ \binom{s-mp+n^{2}-2n}{n^{2}-2n}&\text{else}.\end{cases}$ (21)
where $|S|=(n-2)(p-1)$.
D.3 Empirical results
To empirically determine the solutions to $f(n,p)$ we implemented a brute-force algorithm by iterating over all $p^{(n-1)^{2}}$ candidate matrices of size $(n-1)\times(n-1)$ and verifying whether the two constraints are not violated (see section 4).
The results are given in Figure A9 and Figure A8. Interestingly, $f(n,2)=n!$, but in general $f(n,p)$ scales super-factorially in $n$, for a given $p$.
Figure A9: Number of DSMs of fixed size $n$ with a given number of discretization steps $p$; up to values of $n=10$ and $p=10$. Empty cells require $>5$ days of compute time on a machine with $128$ cores and $128$GB RAM.
Appendix E Quantum hardware experiment
We measured the extent and the effect of quantum hardware noise on a DSM produced by our quantum circuit. To that end, we picked a shallow circuit ($1$ layer) with $14$ qubits and computed the ground-truth $8\times 8$ DSM for a random input matrix through statevector simulation. We then transpiled the circuit on three different quantum computers (Cusco, Brisbane and Torino) available to the public via the IBM Quantum Platform. After using transpilation optimization level $1$, we obtained a circuit with a 2-qubit-depth of $15$ and and a total of $52$ 2 qubit gates. As error mitigation techniques, we used dynamical decoupling [20], Pauli twirling [53] and a projection to the Birkhoff polytope of the approximate-DSM obtained from the quantum circuit (see Section 2.3.2). The results, shown in Figure 6, indicate that, consistently, Cusco was the noisiest machine and Torino yielded the best results. Moreover, in general, beyond a shot count of $10,000$ little performance improvement can be observed. This is a positive finding because it is substantially below the theoretical minimum given by the shot noise limit ($640,000$). However, the deviation from the exact DSM, measured in Frobenius Distance, was substantial (Figure 6A). We analyzed the root cause of this and found that the deviation can be largely attributed to an increase in entropy. DSMs obtained from noisy quantum hardware converge toward the center of the Birkhoff polytope ($1/n$ in every cell). The relative ordering of the absolute values instead is largely preserved (Figure 6C).